# Naturalistic Computational Cognitive Science Towards generalizable models and theories that capture the full range of natural behavior

### v1

@@ 7 Discussion
Conclusions We have outlined a perspective that advocates for studying cognition through a broad spectrum of naturalistic experimental paradigms, building generalizable models of cognition that can perform naturalistic tasks, and deriving explanatory theories from these models and their simplifications. We have supported this perspective with examples illustrating the value of naturalistic settings and generalizable models, and concrete guidance on the practicalities of building generalizable models and using them to shape our theories. Many individual aspects of our argument align with prior works, but we hope that there is value in synthesizing these perspectives and highlighting their synergy. We hope that this perspective will inspire computational cognitive scientists to continue embracing richer naturalistic experimental paradigms and models.
@@ Acknowledgements
Sam Gershman for guidance and feedback on this article and on how to integrate cognitive science and machine learning methodology. We thank Maria Eckstein, Tyler Bonnen, Jay McClelland, Mike Mozer, Rachel Calcott, Nick Blauch, and Shuze Liu for thoughtful comments and discussions. We thank Felix Hill for many inspirational conversations on generalization, cognition, and more.

### v2

@@ 7.2 Conclusions
We have outlined a perspective that advocates for studying cognition through a broad spectrum of naturalistic experimental paradigms, building generalizable models of cognition that can perform naturalistic tasks, and deriving explanatory theories from these models and their simplifications. We have supported this perspective with examples illustrating the value of naturalistic settings and generalizable models, and concrete guidance on the practicalities of building generalizable models and using them to shape our theories. Many individual aspects of our argument align with prior works, but we hope that there is value in synthesizing these perspectives and highlighting their synergy. We hope that this perspective will inspire computational cognitive scientists to continue embracing richer naturalistic experimental paradigms and models.
@@ Acknowledgements
We first thank Sam Gershman for his extensive guidance and feedback on this article and on how to integrate cognitive science and machine learning methodologies. We also thank Maria Eckstein, Tyler Bonnen, Jay McClelland, Mike Mozer, Rachel Calcott, Nick Blauch, Aran Nayebi, Ruben Cohen, Greta Tuckute, Fenil Doshi and Shuze Liu for thoughtful comments and discussions. We are also grateful for this work being supported by a gift from the Chan Zuckerberg Initiative Foundation to establish the Kempner Institute for the Study of Natural and Artificial Intelligence. Finally, we thank Felix Hill for many inspirational conversations on generalization, cognition, and more.